OpenStudy (anonymous):

The equation of a circle is (x + 7)^2 + (y + 2)^2 = 49. Determine the length of the diameter.

OpenStudy (anonymous):

A. 7 B. 49 C. 9 D. 14

OpenStudy (anonymous):

Use the formula of a circle (x - a) ^ 2 + (y - b) ^ 2 = r ^2 r = radius so d = 2 x radius.

OpenStudy (anonymous):

still confused, im not that good at geometry

OpenStudy (anonymous):

r^2 is that 49